5. (NPV and IRR) You work in a company that sells furniture. The company is considering a new marketing campaign. The marketing campaign cost is $1M to be paid immediately. You expect that as a result of the campaign, the company will increase its market share and will generate additional annual cash flows of $150,000 forever, starting 1 year fromnow. a. If your company’s cost of capital (the discount rate) is 10%, should it undertake the marketing campaign? Explain. b. What is the marketing campaign’sIRR? How is this computed using excel formulas?
Answer #1

E15 for A B C D E 1 2 Year Cash Flows 3 0 Present Value - 1000000 150000 4 1- Infinity - 1000000 1500000 500000 Net Present VD14 X A B D 1 2 3 Year 0 1- Infinity Cash Flows - 1000000 150000 Net Present Value Present Value =C3 =C4/10% =SUM(D3:04) 4 5